OpenStudy (anonymous):

I know a way by which i can make i can get a total of 120 by using 0,0,0,0,0, and any one mathematical operator? Do You?

OpenStudy (anonymous):

use common sence

OpenStudy (anonymous):

0+0+0+0+0+120

OpenStudy (thegamer007):

(0!+0!+0!+0!+0!)! =(1+1+1+1+1)! =(4)! =5! is 5*4*3*2*1=120

OpenStudy (thegamer007):

hope it helped